OpenStudy (anonymous):

Given that 15x=5y, what would the the ratio of x to y be in simplest form?

OpenStudy (anonymous):

Just divide both sides by 5: 3x=y or 3x:1y

OpenStudy (mr.math):

The ratio of x to y is \(\frac{x}{y}=\frac{5}{15} \text{ or } \frac{1}{3}\).